OpenStudy (pphalke):

a home has a triangular backyard. the second angle of the triangle is 5 degrees more than the first angle. the third angle is 15 degrees more than three times the first angle. find the angles of the triangular backyard

OpenStudy (anonymous):

# angles algebraically x+5 3x+15 x

OpenStudy (pphalke):

@Cadmus_Cain what do I do from there

OpenStudy (pphalke):

180 degrees

OpenStudy (pphalke):

ok

OpenStudy (pphalke):

@Janu16 u know hw to solve this

OpenStudy (janu16):

ik how to do this

OpenStudy (janu16):

give me 5 minutes plz

OpenStudy (janu16):

let 1st angle =x--------------32 degrees 2nd angle=x+5-----------37 degrees 3rd angle=3x+15-----------111 degrees 5x+20=180 -20 -20 5x=160 /5 /5 x=32 degrees

OpenStudy (janu16):

its 32,37,111